A storage tank at STP contains 26.3 kg of nitrogen (N2).

What is the pressure if an additional 27.7 kg of nitrogen is added without changing the temperature?
Answer #1

Using Ideal gas law:

PV = nRT

Step 1: find the volume of tank

V = nRT/P

n = number of moles of N2 = m/Mw = 26.3 kg/(28*10^-3 kg/mol) = 939.28 moles

R = 8.314

T = 0 C = 273 K

P = 1 atm = 1.01325*10^5 Pa, So

V = 939.28*8.314*273/(1.01325*10^5) = 21.04 m^3

Since temperature is constant, So

Now when 27.7 kg Nitrogen is added, then new number of moles will be:

n1 = m1/Mw = (26.3 + 27.7)/(28*10^-3) = 1928.57 moles

V1 = V, and T1 = T, So

P1 = n1*R*T1/V1

P1 = 1928.57*8.314*273/(21.04)

P1 = 208047.4 Pa = 2.08*10^5 Pa

Please Upvote.